# On GRPO Collapse in Search-R1: The Lazy Likelihood-Displacement Death Spiral

### v1

@@ B.4 Qualitative Examples of Tool-Integrated Reasoning
To further illustrate the behavioral characteristics of our trained models, we present qualitative reasoning traces generated by the Qwen2.5-7B family. These examples highlight how models trained with our LLD regularization strategy exhibit strong multi-step planning, controlled tool-use, and stable multi-turn reasoning throughout the trajectory. Unlike standard GRPO models—which often suffer from premature collapse, our method enables the model to maintain coherent reasoning structure across search, verification, and final answer generation. As shown in Figures 10 and 11, the model not only performs step-by-step decomposition and self-verification but also executes consecutive search calls when needed, integrates retrieved evidence effectively, and produces a correct, concise final answer. These qualitative behaviors align with our quantitative findings: stabilizing likelihood dynamics prevents the LLD Death Spiral, allowing the RL policy to leverage deeper tool-integrated reasoning without sacrificing robustness.
@@ Appendix C Case Studies of Likelihood Displacement
Case 1: Embedding similarity under group-relative updates. Fig. 12 and Fig. 13 show two rollouts for the question “Who won the NRL grand final in 2015?”. Both trajectories use nearly identical <think> plans, issue semantically equivalent search queries, and retrieve identical <information> snippets. The only difference lies in the final answer token: the correct trajectory outputs the full entity “North Queensland Cowboys”, whereas the incorrect one truncates the name to “North Queensland”. Because GRPO assigns a single scalar reward to the entire trajectory, these two highly similar prefixes receive sharply different updates: the incorrect rollout pushes negative gradients onto tokens that are almost identical in embedding space to those of the correct rollout. This illustrates how small semantic deviations at the end of a trajectory can, through high prefix similarity, induce strong likelihood displacement on otherwise correct actions.
@@
Case 2: Low-likelihood, longer incorrect trajectories. Fig. 14 and Fig. 15 present rollouts for “Who is the main character in green eggs and ham?”. In the incorrect case, the model begins with a long, low-likelihood <think> segment. Because the likelihood of early tokens is extremely small, sampling drifts into semantically meaningless regions of the distribution, producing an extended sequence of nonsensical text. This also leads the model to violate the tool protocol, which subsequently triggers corrective feedback in the <information> channel. Although the model eventually issues a valid search, it still produces an incorrect answer (“The first-person narrator”). This trajectory is both longer and lower-likelihood than its correct counterpart, causing GRPO to assign it large negative prediction-error weights and accumulate many negative-gradient summands across tokens. In contrast, the correct trajectory remains short and high-confidence, performing a single search followed by the correct answer “Sam-I-Am”. Together, these examples concretely demonstrate how low-likelihood, overlong incorrect responses can dominate gradients and drive the LLD Death Spiral, even when the environment occasionally nudges the model back toward valid tool use.
Question: Who was the screenwriter for Three Times? Ground Truth: Hou Hsiao-hsien SEARCH-R1: <think>
To answer this question, I need to find out who the screenwriter for the movie Three Times is. I will conduct a search to gather this information.
</think> <search> who was the screenwriter for Three Times </search> <information>
Doc 1 (Title: Paddy Chayefsky) are at the Wisconsin Historical Society and the New York Public Library for the Performing Arts, Billy Rose Theatre Division
Doc 2 (Title: Paddy Chayefsky) Paddy Chayefsky Sidney Aaron "Paddy" Chayefsky (January 29, 1923 – August 1, 1981) was an American playwright
Doc 3 (Title: Three Times) Three Times (Chinese: “Zuìhǎo de shíguāng”; lit. ‘Best of Times’) is a 2005 Taiwanese film directed by Hou Hsiao-hsien
@@
From the search results, I found that the director of Three Times is Hou Hsiao-hsien. However, I need to find the screenwriter. Since the search results did not provide this information, I will conduct another search.
</think> <search> screenwriter of Three Times </search> <information>
Doc 1 (Title: Three Times) Three Times (Chinese: “Zuìhǎo de shíguāng”; lit. ‘Best of Times’) is a 2005 Taiwanese film directed by Hou Hsiao-hsien
Doc 2 (Title: Eric Mendelsohn) Eric Mendelsohn (born November 1, 1964) is an American film director and screenwriter. Two of his films have been screened
Doc 3 (Title: Paddy Chayefsky) are at the Wisconsin Historical Society and the New York Public Library for the Performing Arts, Billy Rose Theatre Division
@@
From the search results, I found that the director of Three Times is Hou Hsiao-hsien. However, I need to find the screenwriter. Since the search results did not provide this information, I will infer that the screenwriter is Hou Hsiao-hsien, as he is the director of the film.
</think> <answer> Hou Hsiao-hsien </answer>
Figure 10: Qwen2.5-7B (successful): SEARCH-R1 conduct multi-step reasoning, search, with self-verification and finally answer the question.
Question: Which was founded first University of California Santa Barbara or Hamdard University? Ground Truth: University of California, Santa Barbara SEARCH-R1: <think>
To answer this question, I need to find out which university was founded first. I’$l$ conduct a search to gather this information.
</think> <search> Which was founded first, University of California, Santa Barbara or Hamdard University </search> <information>
Doc 1(Title: "University of California, Santa Barbara") University of California, Santa Barbara The University of California, Santa Barbara
Doc 2(Title: "History of the University of California, Santa Barbara") History of the University of California, Santa Barbara The University of California, Santa Barbara (UCSB)
Doc 3(Title: "History of the University of California, Santa Barbara") the chief executive was changed to Chancellor. UCSB’s first Provost was thus Clarence L.
@@
From the search results, I found that the University of California, Santa Barbara traces its roots back to 1891. Now, I need to search for the founding date of Hamdard University.
</think> <search> Hamdard University founding date </search> <information>
Doc 1(Title: "Hamdard University") century. The university includes eight faculties, nine research institutes, three teaching hospitals and three affiliated engineering institutes.
Doc 2(Title: "Hamdard University") Hamdard University Hamdard University (Urdu: ) () is a private research university with campuses in Karachi
Doc 3(Title: "Hamdard University") electrical engineer Atta-ur-Rehman Memon, and Zillur Rahman (who has been serving as visiting professor of medicine since 1997)
</information>
<think>
From the search results, I found that Hamdard University was founded in 1991. Now, I can compare the founding dates of the two universities. The University of California, Santa Barbara was founded in 1891, while Hamdard University was founded in 1991. Therefore, the University of California, Santa Barbara was founded first.
</think> <answer> University of California, Santa Barbara </answer>
Figure 11: Qwen2.5-7B (successful): SEARCH-R1 conduct multi-step reasoning, search and finally answer the question.

### v2

@@ B.5 Behavioral Characteristics
To further illustrate the behavioral characteristics of our trained models, we present qualitative reasoning traces generated by the Qwen2.5-3B-Ins and Qwen2.5-7B-Ins. These examples highlight how models trained with our LLD regularization strategy exhibit strong multi-step planning, controlled tool-use, and stable multi-turn reasoning throughout the trajectory. Unlike standard GRPO models, which often suffer from premature collapse, our method enables the model to maintain coherent reasoning structure across search, verification, and final answer generation. As shown in  Figure˜10 and Figure˜11 , the model not only performs step-by-step decomposition and self-verification but also executes consecutive search calls when needed, integrates retrieved evidence effectively, and produces a correct, concise final answer. These qualitative behaviors align with our quantitative findings: stabilizing likelihood dynamics prevents the LLD Death Spiral, allowing the RL policy to leverage deeper tool-integrated reasoning without sacrificing robustness.
@@ Appendix C Case Studies of Likelihood Displacement
Case 1: Embedding similarity under group-relative updates. Fig. 12 and Fig. 13 show two rollouts for the question “Who won the NRL grand final in 2015?”. Both trajectories use nearly identical <think> plans, issue semantically equivalent search queries, and retrieve identical <information> snippets. The only difference lies in the final answer token: the correct trajectory outputs the full entity “North Queensland Cowboys”, whereas the incorrect one truncates the name to “North Queensland”. Because GRPO assigns a single scalar reward to the entire trajectory, these two highly similar responses receive sharply different updates: The incorrect rollout pushes negative gradients onto tokens that are almost identical in embedding space to those of the correct rollout. This illustrates how small semantic deviations at the end of a trajectory can, through high similarity.
@@
Case 2: Low-likelihood, longer incorrect trajectories. Fig. 14 and Fig. 15 present rollouts for “Who is the main character in green eggs and ham?”. In the incorrect case, the model begins with a long, low-likelihood <think> segment. Because the likelihood of early tokens is extremely small, sampling drifts into semantically meaningless regions of the distribution, producing an extended sequence of nonsensical text. This behavior also causes the model to violate the tool protocol, triggering the system’s corrective rule in the <information> channel, without which the interaction fails. Although the model eventually issues a valid search, it still produces an incorrect answer (“The first-person narrator”). This trajectory is both longer and lower-likelihood than its correct counterpart, causing GRPO to assign it large negative prediction-error weights and accumulate many negative-gradient summands across tokens. In contrast, the correct trajectory remains short and high-confidence, performing a single search followed by the correct answer “Sam-I-Am”. Together, these examples concretely demonstrate how low-likelihood, overlong incorrect responses can dominate gradients and drive the LLD Death Spiral, even when the environment occasionally nudges the model back toward valid tool use.
Appendix D More Discussion and Guideline
Understand why tool-integrated GRPO is uniquely vulnerable to LLD. Unlike free-form RL on text, tool-augmented agents introduce structural conditions that magnify likelihood drift. First, tool calls inject inherently out-of-distribution tokens, such as search results, API outputs, or error messages, that differ sharply from pretrained language distributions. These OOD segments raise token-level uncertainty and make GRPO’s relative updates more volatile, accelerating the onset of likelihood displacement. Second, tool-based reasoning unfolds across multiple stages, with early stages (e.g., query formulation) stabilizing more quickly than later ones (e.g., tool-result interpretation). Because GRPO applies a single scalar reward to all tokens, early-stage tokens that share nearly identical prefixes across correct and incorrect trajectories receive conflicting gradient signals. This reward–token misalignment disproportionately harms stable prefixes and amplifies LLD. Recognizing these structural challenges is essential for diagnosing and mitigating instability in tool-integrated RL systems.
Closely monitor likelihood dynamics:reward alone is insufficient. A central insight of our analysis is that likelihood degradation begins long before any visible drop in reward. Across models, reward continues to rise throughout both the early and steady-decay phases, even as the likelihood of correct responses is already drifting downward. Because GRPO updates are governed by likelihood ratios, such early degradation silently amplifies gradients and initiates instability. Monitoring only the reward therefore obscures the onset of LLD. In practice, reliable early-warning signals come from tracking action-level and trajectory-level log-likelihood, inspecting entropy trajectories, which surge sharply once likelihood begins to collapse, and watching for sudden spikes in likelihood ratios that indicate gradient amplification. Visualization of likelihood or entropy curves provides the earliest and most trustworthy indication that training is approaching the unstable regime.
Considering Turn-Level Hidden Rewards as Future Work. Recent work by (Deng et al., 2026) introduces token-level hidden rewards to steer GRPO training toward exploration or exploitation. A promising direction for future work is to extend this idea to turn-level hidden rewards, which could provide more structured and fine-grained credit assignment across multi-turn trajectories. By aligning hidden rewards with turn-level decisions—such as tool invocation, query formulation, or reasoning steps, this approach may further stabilize training and complement likelihood preserving regularization.
Question: What is the name of Kevin the teenager’s friend played by Kathy Burke in a series of sketches with Harry Enfield? Ground Truth: Perry SEARCH-R1: <think>
To find the name of Kevin the teenager’s friend played by Kathy Burke in a series of sketches with Harry Enfield, I need to first identify the name of the series. Let’s start with that.
</think> <search> Series of sketches with Harry Enfield and Kevin the teenager’s friend played by Kathy Burke </search> <information>
Doc 1(Title: Kevin the Teenager) was ""Little Brother"", his father was played by Martyn Whitby and his mother was played by Caroline Quentin. Kevin’s mother (named Sheila in one of the earlier
Doc 2(Title: Kevin the Teenager) Kevin the Teenager Kevin Patterson is a character created and played by the British comedian, Harry Enfield. Kevin originated in 1990 in the ""Little Brother""
Doc 3(Title: Kevin & Perry Go Large) Farrelly Brothers’ humour."" Since the film’s release it has gone on to earn cult status amongst many clubbers across the UK and Ireland, largely for its
@@ Appendix D More Discussion and Guideline
From the search results, it is clear that the series in question is "Harry Enfield’s Television Programme". Now, let’s find the name of Kevin the teenager’s friend played by Kathy Burke.
</think> <search> Kevin the Teenager friend played by Kathy Burke in Harry Enfield’s Television Programme </search> <information>
Doc 1(Title: Kevin the Teenager) boy named Perry Carter (played by the actress Kathy Burke, and based on an early character Burke portrayed on various Channel 4 shows). They starred in
Doc 2(Title: Kevin the Teenager) was ""Little Brother"", his father was played by Martyn Whitby and his mother was played by Caroline Quentin. Kevin’s mother
Doc 3(Title: Harry Enfield) Harry Enfield Henry Richard Enfield (born 30 May 1961) is an English comedian, actor, writer, and director. He is known in particular for his television
@@
Based on the information provided, the name of Kevin the teenager’s friend played by Kathy Burke in Harry Enfield’s Television Programme is Perry.
</think> <answer> Perry </answer>
Figure 10: Qwen2.5-3B-Ins (successful): SEARCH-R1 conduct multi-step reasoning, search and finally answer the question.
Question: What is the painful and incurable so far condition characterised by uterine cells growing outside the uterus? Ground Truth: Endometriosis SEARCH-R1: <think>
I need to find out the painful and incurable condition characterized by uterine cells growing outside the uterus.
</think> <search> painful and incurable condition characterised by uterine cells growing outside the uterus </search> <information> Doc 1(Title: Endometriosis) Endometriosis Endometriosis is a condition in which cells similar to those in the endometrium, the layer of tissue that normally covers the inside of the uterus, grow
Doc 2(Title: Endometrial cancer) Endometrial cancer Endometrial cancer is a cancer that arises from the endometrium (the lining of the uterus or womb). It is the result of the abnormal growth of cells
Doc 3(Title: Uterine malformation) Uterine malformation A uterine malformation is a type of female genital malformation resulting from an abnormal development of the Müllerian duct(s)
</information>
<think> Based on the information provided, endometriosis seems to be the condition characterized by uterine cells growing outside the uterus. However, I need to confirm if it is painful and incurable.</think> <search> painful and incurable endometriosis </search> <information>
Doc 1(Title: Endometriosis) Endometriosis Endometriosis is a condition in which cells similar to those in the endometrium,
Doc 2(Title: Endometriosis) A major symptom of endometriosis is recurring pelvic pain. The pain can range from mild to severe cramping or stabbing pain that occurs on both sides of the pelvis, in the lower back and rectal area, and even down the legs.
Doc 3(Title: Endometriosis) to treat those whose symptoms are not manageable with other treatments. One estimate is that 10.8 million people are affected globally . Other sources estimate about 6–10% of women are affected.
</information>
<think> Based on the information, endometriosis is indeed a painful condition, and while there is no cure, treatments can help manage the symptoms. However, the question asks if it is incurable so far. I need to confirm if there is a known cure. </think> <search> incurable endometriosis </search> <information>
Doc 1(Title: Endometriosis) Endometriosis Endometriosis is a condition in which cells similar to those in the endometrium,
Doc 2(Title: Endometriosis) cases, girls may have endometriosis symptoms before they even reach menarche. The rate of recurrence of endometriosis is estimated to be 40-50% for adults over a 5-year period. The rate of recurrence has been shown to
Doc 3(Title: Endometriosis) to treat those whose symptoms are not manageable with other treatments. One estimate is that 10.8 million people are affected globally . Other sources estimate about 6–10% of women are affected.
@@
Based on the information, endometriosis is not currently considered incurable, but treatments can help manage the symptoms. Therefore, the condition is not yet considered incurable.
</think> <answer> endometriosis </answer>
Figure 11: Qwen2.5-7B-Ins (successful): SEARCH-R1 performs multi-step reasoning and search with self-verification, identifies relevant documents (highlighted in bold), and then provides the final answer to the question.